OpenStudy (anonymous):

Turn into slope intercept form a = 400 - 2b a = b + 100

OpenStudy (ash2326):

Suppose a is the independent variable, we know that slope intercept equation for a line with slope m and y-intercept c is y=mx+c here we have a=400-2b or 2b=400-a divide both sides by 2 b=-a/2+200 here slope is -1/2 and intercept of 200 now you try the second part bring b to one side, making sure that its coefficient is +1 you'll have equation in the form b=ma+c